February 10, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, February 10, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .

Calibration Tasks


  1. Announcements
  2. Collaboration Meeting
  3. Calibration Tasks
    • Spring 2016 Recon Checklist
    • Calibration Challenge update
      • Fixed edge case where run number (and CCDB) was not being correctly loaded for some files
      • Moved to Run 3185 to avoid problems with PS hits
      • Successfully reproduced calibrations when starting from current settings
      • Currently refactoring steps
  4. Current Run
  5. Subdetector Reports
  6. Future Simulations
  7. AOB


Attending: Sean (NU); Will, Curtis, Naomi, Mike S. (CMU); Paul, Simon, Elton, Luke, Mark I., Nathan, Dominick, Brad, Adesh, Alex B., Mahmoud, Beni (JLab); Alyssa (FSU)

  1. Announcements
    • Over the last week, Paul modified the offline monitoring scripts to run over incoming data. These scripts are working and will process any new data as it hits the tape.
    • The RCDB is now being updated with run condition information for the current running period and should be considered the canonical source. Data from 2015 is not in there yet.
    • Sean pointed out that if people have questions or problems with GlueX, discussing/asking about them in sim-recon GitHub issue tracker or HyperNews forum will give better results than directly emailing the experts, and will helping in building knowledge and tracking problems.
    • Next offline monitoring launch is this Friday, so changes must be checked in and plugins fixed by 3PM JLab time on Friday.
  2. Collaboration Meeting agenda is set.
  3. Calibration Tasks
    • Sean mentioned a few items about calibration of simulations. Some improvements require changes to the digitization routines, which need to be redone anyway when we move to HGeant4 In the meantime, we should think of what progress can be made by tuning the available parameters, e.g. mcsmear parameters, resolutions, alignments.
    • Sean gave a short update on the Calibration Challenge.
      • He found an edge case where the run number (and therefore CCDB) was not being correctly loaded for some files, and has worked on a solution with David that is being implemented.
      • Moved to Run 3185 to avoid problems with loading PS hits, and successfully reproduced calibrations when starting from current settings (taking into account limited statistics). Jobs seem sufficient to run over reasonably calibrated data.
      • Still studying challenge settings, to better understand interplay of procedures. Currently refactoring the calibration steps.
  4. Current Run
    • Please review the monitoring plugins/web plots that you want created, and let Paul & Sean know of any changes. An overview of the monitoring plans for this run is Data Monitoring Overview.
    • Sean plans to run calibrations jobs on some full detector runs as they come in, to tune the procedure. The outputs could be useful to calibrators. He will prepare a more detailed plan along with some additional accounting.
    • We had a discussion on identifying "production" runs. Paul (and others?) suggested a common set of RCDB queries (e.g. certain trigger settings, solenoid current, etc.). Elton suggested that a minimum number of events cut and additional flag to exclude bad runs would be useful. It might be nice to have some easy way to apply these searches. Sean will start to compile a list, and talk to Dmitry about the best way to interface with the RCDB.
  5. Subdetector Reports
    • FCAL - Adesh reported that they plan to take data as the solenoid is ramping (tomorrow?) to test possible B-field dependence of the gain.
    • BCAL - Will updated the gains in CCDB based on the fall data with updated integration windows.
    • Tracking - Mike is working on calibrations of the new fADC125 data and updating his millipede-based alignment. The latest data is looking much better behaved. Simon reported that Eugene discovered that the survey and GlueX coordinate systems were shifted by ~5 mm, so the geometry will need to be updated when this is better understood. Some of the solenoid coils and other parts may not be in the correct place, which means that the magnetic field maps will need to be updated.
    • TOF - An HV scan is planned for this run. Beni has completed his TDC calibration, and is waiting for data to see some before/after effects.
    • Start Counter' - The status of a potential paddle-dependent efficiency effect in the nose region was discussed. The problem is in the very tip where the paddles are the smallest. The match criteria might need to be changed here. Mahmoud will look into whether there are hits in nearby paddles that should be matched.
    • TAGH - Nathan reported that he has been analyzing the HV scan data taken in the fall, which covered a range of +-150V. The current settings look close to optimal, they might run a little lower to extend the lifetime of the detector. The resolutions are essentially unchanged, varying ~+-40 ps. The efficiencies are more difficult - for some settings, the timewalk calibrations need to be reevaluated and some simulations may be needed.
    • TAGM - Alex reported that he's working on a new method for alignment, since the reduced efficiencies of some of the fiber bundles means that only two of the four signal channels needed for the planned alignment method are being used. He is working with Hovanes on implementing an online procedure using EPICS scalar data, where the microscope is strobed row-by-row.
    • RF - Paul calibrated the RF with recent data. The jitter is still there, but at a smaller level.
    • Trigger-related issues - Nathan is working on updating his PS skim plugin to use the new trigger data. Alex Somov has promised to release a definition of the words stored in the trigger objects. Sean would also like to use this information to look at random triggers. Elton reported that we are currently taking data with mixed cosmic and BCAL LED data, and are working on a plan to optimize the BCAL LED monitoring and learning how to analyze data taken with multiple triggers.
  6. Future Simulations
    • Mark reported that he built a new version of HDDS that has the plastic target on the start counter's nose removed, and a new version of sim-recon. 10 test jobs are on the farm right now.